Exercise 21.1 Part A Constants Excess electrons are placed on a small lead sphere with a mass of 8.10 g so that its net charge is -3.00x10-9 C Find the number of excess electrons on the sphere Submit Part B How many excess electrons are there per lead atom? The atomic number of lead is 82, and its molar mass is 207 g/mol Wead Submit
Answer #1

A)

No. of electrons, n = q/e = (3 x 10^-9)/(1.6 x 10^-19) = 1.88 x 10^10

B)

N = (8.10/207) x 6.022 x 10^23 = 2.36 x 10^22

N' = (1.88 x 10^10)/(2.36 x 10^22) = 7.98 x 10^-13

Comment in case any doubt please rate my answer ....